OpenStudy (anonymous):

Please help will medal!!! 6+log9 (r−5)=5

OpenStudy (anonymous):

ok so remember log_a (x) = q <=> a^(q)= x change the equasion so that the 6 is at the other side of the = sign. That makes -1 9^(-1)=r-5 What is r?

OpenStudy (anonymous):

46/9